OpenStudy (anonymous):

Pleassssseeee help Find three consecutive even integers such that the difference between three times the larger and the smaller is 52.

hartnn (hartnn):

let the smallest integer be x so the three even integers will be x, x+2 and x+4 did u get this?

OpenStudy (anonymous):

Do you put that into an equation

hartnn (hartnn):

the difference between three times the larger and the smaller is 52. so 3(x+4) - x = 52 did u get this equation?

OpenStudy (anonymous):

Sure

hartnn (hartnn):

can u solve that?

OpenStudy (anonymous):

That equals 20

hartnn (hartnn):

yup, x=20 3 integers are 20,22,24

OpenStudy (anonymous):

How did you find the other two

hartnn (hartnn):

smallest = x = 20 next = x+2 = 22 largest = x+4 = 24

OpenStudy (anonymous):

Thank you sooooo much

hartnn (hartnn):

welcome sooooo much